If 0.224L of H2 gas is formed at cathode, the volume of O2 gas formed at amode under identical condition?[please write the equation of H2 and O2 ].,

Grade:12th pass

1 Answers

Shatakshi sharma
6 years ago
H2 gas is formed at cathode and o2 at anode.. it is electrolysis of waterH2O ---> H2 + 1/2 O2 this means 1 mole H2O gives 1 mole H2 and 1/2 mole O2.According to questionMoles of H2 :Formula : volume in litres ÷ 22.4L = moles => 0.224( given ) ÷ 22.4 Moles = 1/100 1mole h2 will react with 1/2 mole o2Therefore moles of O2 will be 1/100×2 = 1/200Volume of O2 = moles × 22.4 L ---> 1×22.4/200 => Moles of O2 = 0.112 litres